OpenStudy (anonymous):

HELP PLEASE! multiply k+3/4k-2 times (12k^2-3)

jimthompson5910 (jim_thompson5910):

Hint: Factor 4k-2 to get 2(2k-1) and factor 12k^2-3 to get 3(4k^2-1) which factors further to 3(2k+1)(2k-1)

jimthompson5910 (jim_thompson5910):

You should see some common terms that will cancel

OpenStudy (anonymous):

um i need a bit more help with this @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large \frac{k+3}{4k-2} \times (12k^2-3)\] \[\Large \frac{k+3}{4k-2} \times \frac{12k^2-3}{1}\] \[\Large \frac{k+3}{2(2k-1)} \times \frac{3(2k+1)(2k-1)}{1}\] \[\Large \frac{k+3}{2\cancel{(2k-1)}} \times \frac{3(2k+1)\cancel{(2k-1)}}{1}\] See how I'm getting all this?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

but 3(2k+1)/1 is not one of the multiple choice answers

jimthompson5910 (jim_thompson5910):

one moment

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

\[\Large \frac{k+3}{4k-2} \times (12k^2-3)\] \[\Large \frac{k+3}{4k-2} \times \frac{12k^2-3}{1}\] \[\Large \frac{k+3}{2(2k-1)} \times \frac{3(2k+1)(2k-1)}{1}\] \[\Large \frac{k+3}{2\cancel{(2k-1)}} \times \frac{3(2k+1)\cancel{(2k-1)}}{1}\] \[\Large \frac{k+3}{2} \times \frac{3(2k+1)}{1}\] \[\Large \frac{(k+3)*3(2k+1)}{2*1}\] \[\Large \frac{3(k+3)(2k+1)}{2}\] \[\Large \frac{3(2k^2+k+6k+3)}{2}\] \[\Large \frac{3(2k^2+7k+3)}{2}\] \[\Large \frac{6k^2+21k+9}{2}\]

jimthompson5910 (jim_thompson5910):

Hopefully that all makes sense

OpenStudy (anonymous):

here are the multiple choice answers A.)3(k+3)(2k+1)/2 B) 3(k+3)(2k-1)/2 C) 3(k+3)/2 D) 3(k+3)(k-1)

jimthompson5910 (jim_thompson5910):

ah so they left it factored, ok let me fix it

jimthompson5910 (jim_thompson5910):

\[\Large \frac{k+3}{4k-2} \times (12k^2-3)\] \[\Large \frac{k+3}{4k-2} \times \frac{12k^2-3}{1}\] \[\Large \frac{k+3}{2(2k-1)} \times \frac{3(2k+1)(2k-1)}{1}\] \[\Large \frac{k+3}{2\cancel{(2k-1)}} \times \frac{3(2k+1)\cancel{(2k-1)}}{1}\] \[\Large \frac{k+3}{2} \times \frac{3(2k+1)}{1}\] \[\Large \frac{(k+3)*3(2k+1)}{2*1}\] \[\Large \frac{3(k+3)(2k+1)}{2}\]

jimthompson5910 (jim_thompson5910):

So it looks like choice A

OpenStudy (anonymous):

Ok ya that makes much more sense haha sorry i should have given u the choices to begin with

jimthompson5910 (jim_thompson5910):

that's ok, I'm glad it makes more sense now

OpenStudy (anonymous):

thanks again jim

jimthompson5910 (jim_thompson5910):

np